# Elastic tensor-derived properties of composition-dependent disordered refractory binary alloys using DFPT

### v3

@@ 3.5 Nuclear-relaxation contributions to elastic properties
[…]
@@
In the case of NbTa (Fig. 6), the constituent elements Nb and Ta have similar single-bond covalent atomic radii ($r_{\rm cov}$ = 1.47 Å for Nb and 1.46 Å for Ta [98]), and exhibit a small difference in their elastic constants. The relaxation values of the Nb atoms and Ta atoms vary with lattice position for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$. This is due to differences in the surrounding atoms at each lattice site, which is the nature of disordered systems. The average relaxation value of a Nb atom for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$ in case of $\rm{Nb_{0.25}Ta_{0.75}}$, where Nb atom is the alloying element, is lower than that of the host Ta atoms. Similarly, the average relaxation value of the Ta atom for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$ for $\rm{Nb_{0.75}Ta_{0.25}}$, where the Ta atom is alloying element, is lower than that of the host Nb atoms. The distribution of atomic relaxation values for $\rm{Nb_{0.25}Ta_{0.75}}$ and $\rm{Nb_{0.75}Ta_{0.25}}$ are similar for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$, but different for $\rm{Nb_{0.5}Ta_{0.5}}$, with no common trend observed among $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$.
@@
Fig. 7 shows the decompositions of relaxation values in WTa, for which the constituent elements W and Ta have different single-bond covalent atomic radii ($r_{\rm cov}$ = 1.37 Å for W and 1.46 Å for Ta [98]), and significant differences in their elastic constants. As for NbTa, the relaxation values of the W and Ta atoms vary with lattice position for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$ and $G$. Similar to NbTa, the average relaxation value of the W atoms is lower for $\bar{C}_{11}$, $\bar{C}_{44}$ and $G$ in the case of $\rm{W_{0.25}Ta_{0.75}}$ when W is the alloying element, and the average relaxation value of the Ta atoms is lower for $\bar{C}_{11}$, $\bar{C}_{44}$ and $G$ in case of $\rm{W_{0.75}Ta_{0.25}}$ when Ta is the alloying element. However, for $\bar{C}_{12}$ of WTa, the average relaxation value of the W atoms is greater when W is the alloying element ($\rm{W_{0.25}Ta_{0.75}}$) and the average relaxation value of the Ta atoms is greater in $\rm{W_{0.75}Ta_{0.25}}$, where Ta is alloying element.
@@
Figure 6: Relaxation field (values in GPa) constructed as described in text for $\rm{Nb_{0.25}Ta_{0.75}}$, $\rm{Nb_{0.5}Ta_{0.5}}$, and $\rm{Nb_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 7: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Ta_{0.75}}$, $\rm{W_{0.5}Ta_{0.5}}$, and $\rm{W_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$,  (g)-(i): $\bar{C}_{44}$, and (j)-(l): $G$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@ 4.1 Elastic tensor variations with distortions in lattice
Within continuum elasticity theory [73], the ideal elastic tensor of a cubic crystal (including BCC) has only three independent elastic constants due to the lattice symmetry of the crystal. The number of independent elastic constants will increase when the symmetry of the crystal is broken. Alloying introduces distortions in the perfect lattice due to differences in the atomic radii and quantum mechanical interactions between the constituent atoms [68]. This causes lattice distortions and introduces local strain, which hinders dislocation motion, thus increasing the mechanical strength [99, 100]. Hence it is crucial to correlate the effects of distortions on the elastic tensor due to alloying, and develop computational methods for characterizing these effects at the microscopic level.
@@
To understand the effect of distortions on the elastic constants we define two types of distortions, crystal distortions and lattice distortions. Crystal distortions correspond to changes in the lattice parameter in different directions due to symmetry breaking of the BCC crystal. Lattice distortions refer to atomic motion away from ideal lattice sites due to differences in the electronic (quantum mechanical) interactions between the atoms. Computation of the relaxed-ion elastic tensor using DFPT, instead of computing only $C_{11}$, $C_{12}$ and $C_{44}$, makes it possible to capture crystal distortions and lattice distortions in terms of standard deviations in Voigt averages and nuclear relaxation values in the BCC binary alloys. Non-zero standard deviation values in the Voigt-averaged elastic constants signal crystal distortions ${\rm a}_{x}\neq{\rm a}_{y}\neq{\rm a}_{z}$. In the case of A = 50 at.composition in the AB binary alloys, the relaxed structures retain cubic crystal symmetry (${\rm a}_{x}={\rm a}_{y}={\rm a}_{z}$), leading to ${C}_{11}={C}_{22}={C}_{33}$, ${C}_{12}={C}_{23}={C}_{31}$ and ${C}_{44}={C}_{55}={C}_{66}$, and thus a zero standard deviation in the Voigt-averaged values. Nevertheless, although crystal symmetry is preserved in the 50-50 binary alloys, deviations away from the perfect lattice are still possible, and indeed, these can play an important role in determining the mechanical behavior of high-entropy alloys [101].
@@
Lattice distortions in the relaxed structure of the binary alloys lead to unbalanced forces on the crystal under a finite strain as illustrated schematically in Fig. 8. The magnitude and direction of resultant unbalanced forces at each lattice site depends on the atom type and the atomic environment in the lattice. In disordered binary alloys, atom type and atomic environment differ at each lattice site. Nuclear relaxation values, which contribute to the computation of relaxed-ion elastic constants, provides the measure of resultant unbalanced forces in the crystal under finite strain. While standard deviation values for the binary alloys with A = 50 at.is zero, indicating cubic crystal structure, the nuclear relaxation values are non-zero and are maximum for $\bar{C}_{11}$. This indicates that in A = 50 at.binary alloys, even though there are no crystal distortions, there are lattice distortions in the crystal which lead to unbalanced forces in the crystal. For same-group binary alloys, $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$ are observed to be close to zero indicating minute lattice distortions except for $\bar{C}_{44}$ of NbTa. For $\bar{C}_{44}$ of NbTa, the nuclear relaxation values are finite even though the standard deviation values are zero. This indicates that, in NbTa, the unbalanced forces are oriented in a particular direction as only $\bar{C}_{44}$ has a finite contribution, while $\bar{C}_{11}$ and $\bar{C}_{12}$ has near-zero contributions. Even though on the macroscopic scale, no crystal distortions are observed, at the microscopic scale the unbalanced forces are oriented in a particular direction, which reduces the magnitude of relaxed-ion elastic constants. In case of MoW and NbTa, nuclear relaxation values are near zero (except for $\bar{C}_{44}$ of NbTa) for all compositions, while the standard deviation values are finite for A = 25 at.and 50 at.compositions. This may be due to near-zero unbalanced forces on the atoms under strain or random orientation of the forces on the atoms resulting in near-zero overall relaxation values.
@@ 4.2 Effect of alloying on elastic properties
The input structures used in the DFPT rigid-ion and relaxed-ion calculations can impact the accuracy of the final computed elastic tensors, through residual stresses. In principle, it is possible to correct for any initial stresses present in the starting structure [32]. In the present work, we have utilized fully-relaxed input SQS supercell structures in order to avoid the need for ex post facto corrections and to maintain internal consistency in our calculations. The input structures are obtained for each binary and composition by minimizing the forces on each atom, allowing cell shape, dimensions and relative atomic positions on the lattice to vary. The particular computational procedures followed here have been essential to obtaining elastic constants in remarkably good agreement with experiment.
Elastic constants are fundamental probe of the elastic interactions of atoms in a given material and are correlated with mechanical properties. Due to random arrangement of atoms in disordered alloys, the elastic interactions are complex and difficult to quantify. In single crystal disordered alloys, solid solution effects like solid solution strengthening are the major contributing factors to observed changes in mechanical properties due to alloying. Since the seminal works of Fleisher [102] and Labusch [103], solid solution effects have been correlated with differences in atomic radii and elastic moduli. The fundamental idea behind this is to capture the lattice distortions and associated local strains due to alloying [104, 105]. Even though these models have been widely verified for yield strength, they can be used to understand trends in elastic constants as well [106, 107]. For same-group binary alloys (MoW and NbTa), the differences in single-bond covalent atomic radii and differences in shear moduli of the constituent elements are smaller compared to those of different-group elements. Hence same-group alloys showed a linear trend across the entire composition range for $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$, closely following Vegard’s law. However, in different-group binary alloys (MoNb, MoTa, WNb and WTa) the differences in atomic radii and differences in shear moduli of the constituent elements lead to nonlinear trends, with deviations from Vegard’s law.
@@ 4.3 Nuclear relaxation contributions to the elastic tensor
Elastic constants are fundamental probes of the nature of bonding in a given material. Elastic constant heterogeneity provides a direct measure of varying bond strength in alloys. Higher relaxation values lead to lower elastic constant values, as relaxation values are subtracted from the rigid-ion elastic constants. This effect plays a crucial role in understanding the location of a crack’s initial site in disordered alloys, as locations with higher relaxation values will have lower elastic constants, hence higher stress concentration sites, thus serving as crack initiation sites. The methodology adopted here can be easily extended to multi-component systems where the relaxation values are expected to be even higher due to larger lattice distortions [108].
@@ 5 Conclusions
In this work, the elastic tensors of composition-dependent refractory binary alloys based on Mo, Nb, Ta, and W have been computed using density functional perturbation theory in conjunction with SQS supercell modeling of the composition-dependent disordered BCC structures. The use of the response-based DFPT methodology for computing the elastic tensor made it possible to compute both the rigid-ion and relaxed ion elastic tensors and nuclear-relaxation contributions, in addition to elastic tensor-derived properties. The Voigt-averaged elastic constants $\bar{C}_{11},\bar{C}_{12}$, and $\bar{C}_{44}$ derived from the elastic tensors were used to compute the polycrystalline elastic moduli $E$, $B$, and $G$. Relaxed-ion calculations for the elastic constants corresponding to the inclusion of nuclear relaxation effects under uniform strain are consistent with experiment. Differences between the rigid-ion and relaxed-ion values for the elastic constants were found to lie within the standard deviations of their Voigt averages for the 25, 50, and 75 at.alloy compositions, suggesting that nuclear relaxation contributions to the elastic tensor are relatively small. This made it possible to extend rigid-ion elastic tensor and derived elastic properties to additional alloy compositions with no loss in accuracy. The computed rigid-ion elastic constants are in excellent agreement with available experimental data, demonstrating the ability of our methodology to accurately predict key mechanical properties from first-principles theory. Derived mechanical properties of $B$, $G$ and $E$ showed a linear trend in properties across the compositions for same-group binary alloys and a nonlinear trend in different-group binary alloys. Our calculations showed elastic anisotropy behavior in all of the binary alloys except WTa.
The computed relaxation values were shown to be small and the decomposition of the elastic constants into electronic contributions and nuclear relaxation contributions helps to understand the effects of nuclear relaxations on the symmetry of the elastic tensor. The lattice-dependent nuclear relaxation fields provide detailed quantum mechanical information for potential use in data-driven predictive models. Computation of relaxation values made it possible to map the heterogeneity in elastic constants which can play a crucial role in understanding crack initial site in disordered alloys. This demonstrated ability to perform accurate, experimentally-validated DFPT calculations will enable advanced, microscopic modeling of the mechanical properties of compositionally- and chemically-complex alloys.

### v4

@@ 3.5 Nuclear-relaxation contributions to elastic properties
The method employed in computing the nuclear-relaxation contributions to the elastic tensor makes it possible to decompose it into atom-specific contributions, allowing more detailed modeling of the microscopic origins of elastic response in the binary alloys.Similar to elastic anisotropy plots, the elastic property heterogeneity can be plotted as lattice dependent spatial variations of the elastic properties. The force response internal strain tensor $\Lambda_{p\alpha,j}$ characterizes the force response to the atomic displacement of atom $p$ along direction $\alpha$ when the crystal is under strain ($s_{j}$), while the displacement response internal strain tensor $\Gamma_{p\alpha,i}$ characterizes the displacement response to the atomic displacement of the atom $p$ along direction $\alpha$ when the crystal is under strain ($s_{i}$). The product of these intermediate values summed over all directions for a given atom $p$ gives the relaxation value of a given elastic constant or elastic property ($B$, $G$, $E$ and $\nu$) of atom $p$ under macroscopic strain (Eqs. (16) and (2.4)). To illustrate the resulting spatial variation in a relaxed crystal structure, the relaxation effects are displayed as contour plots in Figs. 6-9 for NbTa and WTa, and for MoNb, MoTa, MoW, and WTa in the Appendix. The supercell is represented as a set of $\{0\,0\,1\}$ planes. A supercell with unit cells stacked in $2\times 2\times 2$ fashion has 4 unique $\{0\,0\,1\}$ planes. These are used to construct 2D layers in order to represent the 3D periodic supercell. Each layer contains the atoms of their respective (0 0 1)(0\,0\,1) plane, and their first nearest neighbors (1NN) from the next plane. For example, layer 1 consists of atoms from the plane 1 and atoms from plane 2 (these are the 1NN of the atoms in plane 1) projected onto plane 1. Interpolation between atomic contributions is used to construct the spatially-dependent relaxation field [ye2018atomic] for each elastic tensor quantity, using a linear function. The color bar in the contour plot has a range varying from minimum (blue) to maximum (yellow) relaxation value.
@@
In the case of NbTa (Fig. 6 & Fig. 7), the constituent elements Nb and Ta have similar single-bond covalent atomic radii ($r_{\rm cov}$ = 1.47 Å for Nb and 1.46 Å for Ta [pyykko2009molecular]), and exhibit a small difference in their elastic properties. The relaxation values of the Nb atoms and Ta atoms vary with lattice position for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$, $E$ and $\nu$. This is due to differences in the surrounding atoms at each lattice site, which is the nature of disordered systems. The average relaxation value of a Nb atom for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$ and $E$ in case of $\rm{Nb_{0.25}Ta_{0.75}}$, where Nb atom is the alloying element, is lower than that of the host Ta atoms while, an inverse behaviour is observed in case of $\nu$. Similarly, the average relaxation value of the Ta atom for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$ and $E$ for $\rm{Nb_{0.75}Ta_{0.25}}$, where the Ta atom is alloying element, is lower than that of the host Nb atoms. The distribution of atomic relaxation values for NbTa are similar for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$ and $E$ with varying magnitudes except for $\nu$. For $\nu$ the behaviour is observed to be opposite to other elastic properties.
@@
Fig. 8 &  9 shows the decompositions of relaxation values in WTa, for which the constituent elements W and Ta have different single-bond covalent atomic radii ($r_{\rm cov}$ = 1.37 Å for W and 1.46 Å for Ta [pyykko2009molecular]), and significant differences in their elastic constants. As for WTa, the relaxation values of the W and Ta atoms vary with lattice position for $\bar{C}_{11}$, $\bar{C}_{12}$, $\bar{C}_{44}$, $B$, $G$, $E$ and $\nu$. Similar to NbTa, the average relaxation value of the W atoms is lower for $\bar{C}_{11}$, $\bar{C}_{44}$, $B$, $G$, $E$ and $\nu$ in the case of $\rm{W_{0.25}Ta_{0.75}}$ when W is the alloying element, and the average relaxation value of the Ta atoms is lower for $\bar{C}_{11}$, $\bar{C}_{44}$, $B$, $G$, $E$ and $\nu$ in case of $\rm{W_{0.75}Ta_{0.25}}$ when Ta is the alloying element. However, for $\bar{C}_{12}$ of WTa, the average relaxation value of the W atoms is greater when W is the alloying element ($\rm{W_{0.25}Ta_{0.75}}$) and the average relaxation value of the Ta atoms is greater in $\rm{W_{0.75}Ta_{0.25}}$, where Ta is alloying element.
@@
Figure 6: Relaxation field (values in GPa) constructed as described in text for $\rm{Nb_{0.25}Ta_{0.75}}$, $\rm{Nb_{0.5}Ta_{0.5}}$, and $\rm{Nb_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@
Figure 7: Relaxation field (values in GPa) constructed as described in text for $\rm{Nb_{0.25}Ta_{0.75}}$, $\rm{Nb_{0.5}Ta_{0.5}}$, and $\rm{Nb_{0.75}Ta_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 8: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Ta_{0.75}}$, $\rm{W_{0.5}Ta_{0.5}}$, and $\rm{W_{0.75}Ta_{0.25}}$. (a)-(c): $\bar{C}_{11}$,  (d)-(f): $\bar{C}_{12}$ and (g)-(i): $\bar{C}_{44}$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
Figure 9: Relaxation field (values in GPa) constructed as described in text for $\rm{W_{0.25}Ta_{0.75}}$, $\rm{W_{0.5}Ta_{0.5}}$, and $\rm{W_{0.75}Ta_{0.25}}$. (a)-(c): $B$,  (d)-(f): $G$,  (g)-(i): $E$ and (j)-(l): $\nu$. The color bar is set to range from minimum to maximum relaxation values for all atoms. $x$ and $y$ coordinates correspond to reduced coordinates in the 3D supercell.
@@ 4.1 Elastic tensor variations with distortions in lattice
Within continuum elasticity theory [zhao2007first], the ideal elastic tensor of a cubic crystal (including BCC) has only three independent elastic constants due to the lattice symmetry of the crystal. The number of independent elastic constants will increase when the symmetry of the crystal is broken. Alloying introduces distortions in the perfect lattice due to differences in the atomic radii and quantum mechanical interactions between the constituent atoms [bijjala2025composition]. This causes lattice distortions and introduces local strain, which hinders dislocation motion, thus increasing the mechanical strength [senkov2010refractory, zhang2014microstructures]. Hence it is crucial to correlate the effects of distortions on the elastic tensor due to alloying, and develop computational methods for characterizing these effects at the microscopic level.
@@
To understand the effect of distortions on the elastic constants we define two types of distortions, crystal distortions and lattice distortions. Crystal distortions correspond to changes in the lattice parameter in different directions due to symmetry breaking of the BCC crystal. Lattice distortions refer to atomic motion away from ideal lattice sites due to differences in the electronic (quantum mechanical) interactions between the atoms. Computation of the relaxed-ion elastic tensor using DFPT, instead of computing only $C_{11}$, $C_{12}$ and $C_{44}$, makes it possible to capture crystal distortions and lattice distortions in terms of standard deviations in Voigt averages and nuclear relaxation values in the BCC binary alloys. Non-zero standard deviation values in the Voigt-averaged elastic constants signal crystal distortions ${\rm a}_{x}\neq{\rm a}_{y}\neq{\rm a}_{z}$. In the case of A = 50 at.% composition in the AB binary alloys, the relaxed structures retain cubic crystal symmetry (${\rm a}_{x}={\rm a}_{y}={\rm a}_{z}$), leading to ${C}_{11}={C}_{22}={C}_{33}$, ${C}_{12}={C}_{23}={C}_{31}$ and ${C}_{44}={C}_{55}={C}_{66}$, and thus a zero standard deviation in the Voigt-averaged values. Nevertheless, although crystal symmetry is preserved in the 50-50 binary alloys, deviations away from the perfect lattice are still possible, and indeed, these can play an important role in determining the mechanical behavior of high-entropy alloys [owen2018lattice].
@@
Lattice distortions in the relaxed structure of the binary alloys lead to unbalanced forces on the crystal under a finite strain as illustrated schematically in Fig. 10. The magnitude and direction of resultant unbalanced forces at each lattice site depends on the atom type and the atomic environment in the lattice. In disordered binary alloys, atom type and atomic environment differ at each lattice site. Nuclear relaxation values, which contribute to the computation of relaxed-ion elastic constants, provides the measure of resultant unbalanced forces in the crystal under finite strain. While standard deviation values for the binary alloys with A = 50 at.% is zero, indicating cubic crystal structure, the nuclear relaxation values are non-zero and are maximum for $\bar{C}_{11}$. This indicates that in A = 50 at.% binary alloys, even though there are no crystal distortions, there are lattice distortions in the crystal which lead to unbalanced forces in the crystal. For same-group binary alloys, $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$ are observed to be close to zero indicating minute lattice distortions except for $\bar{C}_{44}$ of NbTa. For $\bar{C}_{44}$ of NbTa, the nuclear relaxation values are finite even though the standard deviation values are zero. This indicates that, in NbTa, the unbalanced forces are oriented in a particular direction as only $\bar{C}_{44}$ has a finite contribution, while $\bar{C}_{11}$ and $\bar{C}_{12}$ has near-zero contributions. Even though on the macroscopic scale, no crystal distortions are observed, at the microscopic scale the unbalanced forces are oriented in a particular direction, which reduces the magnitude of relaxed-ion elastic constants. In case of MoW and NbTa, nuclear relaxation values are near zero (except for $\bar{C}_{44}$ of NbTa) for all compositions, while the standard deviation values are finite for A = 25 at.% and 50 at.% compositions. This may be due to near-zero unbalanced forces on the atoms under strain or random orientation of the forces on the atoms resulting in near-zero overall relaxation values.
@@ 4.2 Effect of alloying on elastic properties
The input structures used in the DFPT rigid-ion and relaxed-ion calculations can impact the accuracy of the final computed elastic tensors, through residual stresses. In principle, it is possible to correct for any initial stresses present in the starting structure [romero2020abinit]. In the present work, we have utilized fully-relaxed input SQS supercell structures in order to avoid the need for ex post facto corrections and to maintain internal consistency in our calculations. The input structures are obtained for each binary and composition by minimizing the forces on each atom, allowing cell shape, dimensions and relative atomic positions on the lattice to vary. The particular computational procedures followed here have been essential to obtaining elastic constants in remarkably good agreement with experiment.
Elastic constants are fundamental probe of the elastic interactions of atoms in a given material and are correlated with mechanical properties. Due to random arrangement of atoms in disordered alloys, the elastic interactions are complex and difficult to quantify. In single crystal disordered alloys, solid solution effects like solid solution strengthening are the major contributing factors to observed changes in mechanical properties due to alloying. Since the seminal works of Fleisher [fleischer1963substitutional] and Labusch [labusch1970statistical], solid solution effects have been correlated with differences in atomic radii and elastic moduli. The fundamental idea behind this is to capture the lattice distortions and associated local strains due to alloying [he2018lattice, toda2015modelling]. Even though these models have been widely verified for yield strength, they can be used to understand trends in elastic constants as well [tian2015elastic, delczeg2009ab]. For same-group binary alloys (MoW and NbTa), the differences in single-bond covalent atomic radii and differences in shear moduli of the constituent elements are smaller compared to those of different-group elements. Hence same-group alloys showed a linear trend across the entire composition range for $\bar{C}_{11}$, $\bar{C}_{12}$ and $\bar{C}_{44}$, closely following Vegard’s law. However, in different-group binary alloys (MoNb, MoTa, WNb and WTa) the differences in atomic radii and differences in shear moduli of the constituent elements lead to nonlinear trends, with deviations from Vegard’s law.
@@ 4.3 Nuclear relaxation contributions to the elastic tensor
Elastic constants are fundamental probes of the nature of bonding in a given material. Elastic constant heterogeneity provides a direct measure of varying bond strength in alloys. Higher relaxation values lead to lower elastic constant values, as relaxation values are subtracted from the rigid-ion elastic constants. This effect plays a crucial role in understanding the location of a crack’s initial site in disordered alloys, as locations with higher relaxation values will have lower elastic constants, hence higher stress concentration sites, thus serving as crack initiation sites. The methodology demonstrated here can be easily extended to multi-component systems where the relaxation values are expected to be even higher due to larger lattice distortions [miracle2017critical].
@@ 5 Conclusions
In this work, the elastic tensors of composition-dependent refractory binary alloys based on Mo, Nb, Ta, and W have been computed using density functional perturbation theory in conjunction with SQS supercell modeling of the composition-dependent disordered BCC structures. The use of the response-based DFPT methodology for computing the elastic tensor made it possible to compute both the rigid-ion and relaxed ion elastic tensors and nuclear-relaxation contributions, in addition to elastic tensor-derived properties. The Voigt-averaged elastic constants $\bar{C}_{11},\bar{C}_{12}$, and $\bar{C}_{44}$ derived from the elastic tensors were used to compute the polycrystalline elastic moduli $E$, $B$, and $G$. Relaxed-ion calculations for the elastic constants corresponding to the inclusion of nuclear relaxation effects under uniform strain are consistent with experiment. Differences between the rigid-ion and relaxed-ion values for the elastic constants were found to lie within the standard deviations of their Voigt averages for the 25, 50, and 75 at.% alloy compositions, suggesting that nuclear relaxation contributions to the elastic tensor are relatively small. This made it possible to extend rigid-ion elastic tensor and derived elastic properties to additional alloy compositions with no loss in accuracy. The computed rigid-ion elastic constants are in excellent agreement with available experimental data, demonstrating the ability of our methodology to accurately predict key mechanical properties from first-principles theory. Derived mechanical properties of $B$, $G$, and $E$ showed a linear trend across the compositions for same-group binary alloys and a nonlinear trend in different-group binary alloys. Our calculations showed elastic anisotropy behavior in all of the binary alloys except WTa.
The computed relaxation values were shown to be small and the decomposition of the elastic constants into electronic contributions and nuclear relaxation contributions helps to understand the effects of nuclear relaxations on the symmetry of the elastic tensor. The lattice-dependent nuclear relaxation fields provide detailed quantum mechanical information for potential use in data-driven predictive models of more complex alloys (ternary and beyond). Such models will be essential in order to overcome computational bottlenecks and scaling issues in property-driven materials design as system sizes grow and compositional complexity increases. Computation of relaxation values made it possible to map the heterogeneity in elastic constants which can play a crucial role in understanding crack initial site in disordered alloys. This demonstrated ability to perform accurate, experimentally-validated DFPT calculations will enable advanced, microscopic modeling of the mechanical properties of compositionally- and chemically-complex alloys.